Project

General

Profile

NDOS Commissioning Meeting 2012/02/27

Attendance: Jon, Gavin, Mat, Enrique, Kurt, Carl, Philip, Minerba, Jarek, Susan, Nick, Jan,

Agenda:

  1. Beam and Detector Update
  2. Installation Plans for the week
  3. DAQ Update and Plans for the week
  4. DCS Update and Plans for the week
  5. AOB

Beam and Detector Report

  • Nothing major to report.

Installation plans for this week

  • APD gas drying system (Mat): tests last week on diffusion model showed poor results. Switched over to flow-thru, which worked ok, but new parts are required. No idea when parts will be ready.

DAQ Update and Plans for next week

  • Kurt: possibly roll out recent changes to SW pertaining to partitioning.

DCS Update and Plans for next week

  • No report

DataCheck update

AOB

  • None